OpenStudy (anonymous):

Is knowing the coordinates of the vertex of a parabola enough to determine the domain and range? If so, how? Thank you!

OpenStudy (anonymous):

as long as you know which direction it opens

OpenStudy (anonymous):

Would that be determined if knowing whether the vertex is a maximum or a minimum?

Directrix (directrix):

Let's practice. Give the domain and range of a parabola with vertex (1, -4). @Catseyeglint911 @lonnie455rich And, the answer is ...?

OpenStudy (anonymous):

To be honest, I have no idea how to answer that.

OpenStudy (isaiah.feynman):

The vertex is only useful in finding range. The y co ordinate of the vertex is used.

OpenStudy (anonymous):

if we assume that it opens up or down we know that the domain is \[(-\infty,\infty)\] and if it opens up the range would be \[[-4,\infty)\]

OpenStudy (anonymous):

and yes if the vertex is a maximum it opens down if it is a minimum it opens up

OpenStudy (anonymous):

Thank you all so very much!! :D

OpenStudy (anonymous):

if you graph it, it should be painfully obvious.

OpenStudy (anonymous):

I unfortunately do not have the ability to graph currently.

OpenStudy (anonymous):

do you not have a piece of paper?

Directrix (directrix):

This --> Give the domain and range of a parabola with vertex (1, -4). has no answer without additional information. @Catseyeglint911 Although, I am now thinking that for all parabolas that are functions, the domain will always be the set of Real numbers. The Range is the tricky part.

OpenStudy (anonymous):

yep, the answer is useless without assumptions or special conditions

Directrix (directrix):

@Isaiah.Feynman This message from the calculator link you posted gives me pause in that there is no description of the software to be downloaded.

OpenStudy (anonymous):

I just use wolfram alpha if my calculator isn't around me

Directrix (directrix):

Is this question answerable: Give the domain and range of a parabola with vertex (1, -4) where the vertex is a maximum of the function.

OpenStudy (anonymous):

yes because we know that it opens down

Directrix (directrix):

Domain is the set of Reals. Range is what? @lonnie455rich

Directrix (directrix):

@Catseyeglint911 I think that your question as originally posted has been answered in this thread. Do you agree? If so, what is the answer to your posted question?

OpenStudy (anonymous):

\[(-\infty,-4]\] is the range